OpenStudy (sugamel1):

A compound has carbon, hydrogen and sulfur and after burning, results in 6.601 grams of CO2, 5.406 grams of water and 9.615 grams of SO2. a) Find the empirical formula b) Mass of the sample before burning

OpenStudy (teopisteo):

M(CO2)=12,01+32,00=44,01 g/mol; n(CO2)=m/M= 0,15 mol= n(C) M(H2O)=2,016+16=18,016 g/mol; n(H2O)=m/M= 0,3 mol=> n(H)=0,6 M(SO2)=32,06+32=64,06 g/mol; n(SO2)=m/M= 0,15 mol= n(S) n(C):n(H):n(S)= 1:4:1 => CH4S M(CH4S)=12+4,032+32,06= 48,092 g/mol => m(CH4S)=n.M= 7,214g